OpenStudy (anonymous):

Find the 7th term of the sequence. -1,-2,-4,-8,.....

OpenStudy (anonymous):

do you know what the sequence is?

Parth (parthkohli):

\(T_n = r^{n - 1}\times a_1 \)

OpenStudy (anonymous):

This is an Geometric progression where: first term is -1 and common ratio is : -2/(-1) = 2

Parth (parthkohli):

What is \(a_1\) and the ratio here?

OpenStudy (anonymous):

So what is the 7th term of the sequence?

OpenStudy (anonymous):

\[\large t_7 = (a) \cdot r^{7-1}\] Put a and r here...

OpenStudy (shane_b):

The equation to find the nth term would be \[−(2^{n−1})\] ...just plug in 7 for n.

OpenStudy (anonymous):

@lovejones2012 do you see what each number is multiplied by?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

it's multiplied by 2 -1 x 2=-2 x 2=-4 x 2=-8 x 2=?

OpenStudy (anonymous):

@lovejones2012 do you understand now?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

everything is multiplied by 2